# The Spectral Edge Thesis: A Mathematical Framework for Intra-Signal Phase Transitions in Neural Network Training

### 30 Mar 2026 v1

@@ Abstract
Empirical summary. Gap dynamics in the rolling-window Gram matrix precede every grokking event observed: 24/24 runs with weight decay, 0/24 without. Across six model families spanning 150K–124M parameters (TinyStories 51M, GPT-2 124M, and grokking experiments on Dyck-1, SCAN, and modular arithmetic), 19 of 20 quantitative predictions of the framework are confirmed (7/8 in this paper; the remainder in companion papers [33, 34, 37, 35, 36]). The number of simultaneously active modes is empirically small ($k^{*}\leq 3$) and optimizer-dependent: Muon drives $k^{*}=1$ while AdamW produces $k^{*}=2$ on the same TinyStories 51M model, both reaching comparable validation loss.
@@ 1.2 The Aspect Ratio Regime
The aspect ratio is $\gamma=p/W$. In our setting, $\gamma\sim 10^{4}$$10^{8}$. We call this the extreme aspect ratio regime: $\gamma\to\infty$ with $W$ fixed.
@@
Classical random matrix theory studies $p,n\to\infty$ with $p/n\to\gamma\in(0,\infty)$ fixed. Our regime is $W$ fixed, $p\to\infty$. This simplifies the theory: all $W$ eigenvalues of $\bm{G}$ are observable, and by the law of large numbers, noise eigenvalues (if present) concentrate tightly. The fluctuation theory (Tracy–Widom) gives corrections—but as we shall show, the noise is negligible in practice, and the relevant structure is within the signal.
@@ 2 The Three Axioms
[…]
Axiom 2.2 is justified by: (i) the Hessian spectrum of neural networks concentrates in $O(k)$ large eigenvalues with a near-zero bulk (Sagun et al. [21], Ghorbani et al. [16]); (ii) gradient descent projects onto the top Hessian eigendirections (Gur-Ari et al. [17]); and (iii) empirically, the maximum ratio $d_{k^{*}}/d_{k^{*}+1}$ peaks at $k^{*}=2$ (TinyStories, ratio $\sim 1.79$) and $k^{*}=3$ (GPT-2, ratio $\sim 1.12$).
@@ 3 The Central Question
The spectral edge framework is architecture-agnostic. The Gram matrix, signal flow ODE, gap dynamics, stability coefficient $\alpha_{j}$, loss decomposition, evolving-NTK signal equation, adiabatic parameter $\mathcal{A}$, and the edge-of-stability identification all hold for any differentiable model with a well-defined NTK—MLPs, CNNs, transformers, or otherwise.
@@ 4 The Noise is Negligible
We first establish that the classical BBP framework is vacuous in the extreme aspect ratio regime.
@@ 4.1 The BBP Detection Threshold
$\boxed{\frac{\sigma_{W}}{d_{\mathrm{crit}}}\;\sim\;20\text{--}300.}$ (6)
[…]
@@ 6.3 Null Distribution of the Maximum Ratio
Under the null hypothesis $H_{0}$: all signal strengths are equal ($d_{1}=d_{2}=\cdots=d_{W}$), the ordered eigenvalues of the $W\times W$ Gram matrix follow a Laguerre ensemble. The distribution of the maximum consecutive ratio $R_{\max}=\max_{j}\lambda_{j}/\lambda_{j+1}$ satisfies, for $W$ fixed and $p\to\infty$:
@@ 9 Subspace Stability: The Davis–Kahan Connection
The key observation is that the Davis–Kahan $\sin\Theta$ theorem does not care what lies below the gap—noise or subdominant signal. It only cares about the size of the gap.
@@ 10.1 Dyson Eigenvalue Dynamics
Corollary 10.2 (Eigenvalue Repulsion).
@@ 12.3 Simplified Signal Flow
Weight decay must sit between the NTK eigenvalues of the generalizing and memorization modes. Too small: memorization persists (no gap opens). Too large: generalization is also suppressed. When $\omega=0$: all low-curvature modes persist at $d_{j}\propto 1/\sqrt{h_{j}}\to\infty$, the generalizing direction is buried, and grokking never occurs. This is consistent with the universal finding of 24/24 grokking with $\omega>0$ and 0/24 without (Section 27).
@@ 17.1 Why the Largest Gap: The Gap Maximality Principle
The gap maximality principle is confirmed by every phase transition observed in our experiments (Section 21Section 28): all 24/24 grokking events with weight decay, the $k^{*}$ shifts in GPT-2 and TinyStories, and the Dyck-1/SCAN gap openings all occur at the position of the largest consecutive eigenvalue ratio, never at secondary gaps.
@@ 27 Test 7: Gap Opening and Grokking (Dyck/SCAN)
Proposition 10.7 predicts that gap opening corresponds to capability gain, and Remark 10.8 maps grokking onto a delayed gap opening event within the signal hierarchy. We test this in a setting where the capability gain is unambiguous: the grokking transition, where generalization accuracy jumps from near-zero to near-perfect.
@@
We analyze weight-matrix singular values $\sigma_{1}\geq\sigma_{2}\geq\cdots$ of the query projection $W_{Q}$ during training of 2-layer transformers on Dyck-1 language ($d_{\mathrm{model}}=128$, $\sim$}150K parameters) and 6-layer encoder–decoder transformers on SCAN compositional generalization ($d_{\mathrm{model}}=256$, $\sim$}1.5M parameters). Each task is trained with weight decay $\lambda\in\{0,1.0\}$ across 3 seeds, giving 12 runs total. The gap ratio $R(t)=\sigma_{1}(t)/\sigma_{2}(t)$ is computed at each checkpoint. Grokking is defined as the first step where test accuracy exceeds 0.95.
@@ Results.
    Controls: Without weight decay, no gap opens and no grokking occurs—6/6 grokking runs with $\lambda=1.0$, 0/6 with $\lambda=0$ (100% hit rate, 0% false positives).
@@
Figure 4: Grokking as a spectral edge event (Dyck-1). Grokking runs (weight decay $\lambda=1.0$, red/solid, 3 seeds) vs. control runs ($\lambda=0$, blue/dashed, 3 seeds). Top: singular value ratio $\sigma_{1}/\sigma_{2}$ of $W_{Q}$. With weight decay, the ratio rises dramatically as $\sigma_{2}\to 0$ (gap opens); without weight decay, the ratio stays flat (no gap). Middle: Frobenius norm $\|W_{Q}\|_{F}$. Weight decay compresses the weights; the control runs retain large norms. Bottom: test accuracy. Grokking (accuracy $\to 1$) occurs only in the weight-decay runs, coinciding with the gap opening above. All 6 grokking runs show gap opening; all 6 controls show neither gap opening nor grokking.
@@
  • Dyck/SCAN: gap opens at grokking (Proposition 10.7), with $R$ rising from ${\sim}1.1$ to $10^{3}$.
  • Modular arithmetic: sub-leading gap closes before grokking, triggering a $k^{*}$ shift from position 2–3 to position 1. This is a gap closing event (at the old $k^{*}$) followed by a gap opening event (at the new $k^{*}=1$).
Both mechanisms are predicted by the gap dynamics (Theorem 14.2): gap closing destabilizes the subspace, and gap opening at a new position stabilizes the generalizing direction. The weight decay $\to$ grokking correspondence is universal: 24/24 grok with $\lambda>0$, 0/24 without, across all four task families (48 runs total).
@@
Weight decay plays a dual role across all experiments: in Dyck/SCAN, it drives subdominant singular values toward zero (opening the gap); in modular arithmetic, it compresses the sub-leading eigenvalue spectrum (closing $g_{23}$), which triggers spectral symmetry-breaking and a $k^{*}$ shift. The 100% correspondence between weight decay and grokking across 48 runs (24 grok, 24 control) spanning four task families is consistent with weight decay being a key driver by which gap dynamics are associated with the generalization transition.
@@
The grokking experiments analyze weight-matrix spectra ($\sigma_{j}(W_{Q})$ for Dyck/SCAN, eigenvalues $\lambda_{j}$ for modular arithmetic) rather than the trajectory Gram spectrum $\sigma_{j}(\bm{G})$. The Davis–Kahan stability framework (Remark 16.3) applies to both: gap dynamics control subspace stability regardless of the matrix being analyzed. The universality of the gap $\to$ capability correspondence across both spectral types provides strong evidence for the generality of the framework.
@@ 29 Summary of Theory–Experiment Match
Gap dynamics (10.7) gap event $\Rightarrow$ capability 24/24 grok, 0/24 ctrl
@@ 40 Connection to the Edge of Stability
Across all models studied ($k^{*}\in\{2,3\}$ for GPT-2 124M and TinyStories), the number of simultaneously active modes is small. A plausible explanation is that when $k>k_{\mathrm{crit}}$ modes simultaneously overshoot, nonlinear interactions destabilise the self-correction mechanism of GD, so the system self-organises to keep only a few modes at the edge. This would make $k^{*}$ a property of the optimizer rather than the task or architecture. We do not have a proof of this; it remains an open question.

### 03 Apr 2026 v2

@@ Abstract
Empirical summary. Gap dynamics in the rolling-window Gram matrix precede every grokking event observed: 24/24 runs with weight decay, 1/24 without. Across six model families spanning 150K–124M parameters (TinyStories 51M, GPT-2 124M, Dyck-1, SCAN, and modular arithmetic single- and multi-task), 19 of 20 quantitative predictions of the framework are confirmed (7/8 in this paper; the remainder in companion papers [32, 33, 34, 37, 35, 36]). The number of simultaneously active modes is empirically small ($k^{*}\leq 3$) and optimizer-dependent: Muon drives $k^{*}=1$ while AdamW produces $k^{*}=2$ on the same TinyStories 51M model, both reaching comparable validation loss.
@@ 1.2 The Aspect Ratio Regime
The aspect ratio is $\gamma=p/W$. In our setting, $\gamma\sim 10^{4}$$10^{8}$. We call this the extreme aspect ratio regime: $\gamma\to\infty$ with $W$ fixed. [^1] The symbol $\gamma$ is reused in later sections for eigenvalue gaps ($\gamma_{k}=\lambda_{k}-\lambda_{k+1}$, LABEL:sec:repulsion) and off-diagonal couplings (LABEL:sec:full-gap-flow); context and subscripts disambiguate.
@@
Classical random matrix theory [4, 6] studies $p,n\to\infty$ with $p/n\to\gamma\in(0,\infty)$ fixed. Our regime is $W$ fixed, $p\to\infty$. This simplifies the theory: all $W$ eigenvalues of $\bm{G}$ are observable, and by the law of large numbers, noise eigenvalues (if present) concentrate tightly. The fluctuation theory (Tracy–Widom) gives corrections—but as we shall show, the noise is negligible in practice, and the relevant structure is within the signal.
@@ 2 The Three Axioms
Axiom 2.1 is justified by the empirical observation that in the extreme aspect ratio regime ($p\sim 10^{8}$, $W\sim 10$), the BBP detection threshold $d_{\mathrm{crit}}=\nu(p(W-1))^{1/4}$ evaluates to $\sim 0.2$$1.5$, while the smallest observed singular value $\sigma_{W}\sim 12$$80$. The ratio $\sigma_{W}/d_{\mathrm{crit}}\sim 9$$300$ (see Section 4). Every eigenvalue is at least an order of magnitude above the noise detection threshold.
Axiom 2.2 is justified by: (i) the Hessian spectrum of neural networks concentrates in $O(k)$ large eigenvalues with a near-zero bulk (Sagun et al. [21], Ghorbani et al. [16], Martin and Mahoney [18]); (ii) gradient descent projects onto the top Hessian eigendirections (Gur-Ari et al. [17]); and (iii) empirically, the maximum ratio $d_{k^{*}}/d_{k^{*}+1}$ peaks at $k^{*}=2$ (TinyStories, ratio $\sim 1.79$) and $k^{*}=3$ (GPT-2, ratio $\sim 1.12$).
@@ 3 The Central Question
The spectral edge framework is architecture-agnostic. The Gram matrix, signal flow ODE, gap dynamics, stability coefficient $\alpha_{j}$, loss decomposition, evolving-NTK signal equation, adiabatic parameter $\mathcal{A}$, and the edge-of-stability identification all hold for any differentiable model with a well-defined NTK [22]—MLPs, CNNs, transformers, or otherwise.
@@ 4 The Noise is Negligible
We first establish that the classical BBP framework [1, 2, 3] is vacuous in the extreme aspect ratio regime.
@@ 4.1 The BBP Detection Threshold
$\boxed{\frac{\sigma_{W}}{d_{\mathrm{crit}}}\;\gtrsim\;9\text{--}300.}$ (6)
Every eigenvalue of $\bm{G}$ is at least an order of magnitude above the BBP detection threshold. The BBP phase transition is never approached.
@@ 6.1 Definition of $k^{*}$
Remark 6.3 (Argmax vs. dynamical spectral edge).
The argmax definition (Definition 6.1) identifies the settled gap—the pair of modes already well separated. Empirically, the ratio with the strongest cross-correlation to validation loss sits one position further out ($k^{*}_{\mathrm{dyn}}=k^{*}_{\mathrm{argmax}}+1$), because that is the active frontier: the mode with the smallest protective gap rotates fastest (Davis–Kahan) and tracks training dynamics most tightly (Remark 22.1). When the body text refers to $k^{*}=2$ for TinyStories or $k^{*}=3$ for GPT-2, it uses the dynamical (cross-correlation) convention; the corresponding argmax values are $k^{*}=1$ and $k^{*}=2$. Both conventions satisfy $k^{*}\ll W$.
Remark 6.4 (Signal-weighted $k^{*}$ for moderate aspect ratios).
In the extreme aspect ratio regime ($p/W\sim 10^{7}$), the bare argmax in Definition 6.1 is well-behaved: all singular values are large and trailing ratios are ${\sim}1$. At moderate aspect ratios ($p/W\sim 10^{4}$), trailing singular values can be small enough that ratios between near-zero eigenvalues ($\sigma_{9}/\sigma_{10}=0.28/0.14=2.0$) dominate the argmax, yielding spurious $k^{*}$ values of 8899. A signal-weighted variant suppresses this artifact:
$k^{*}_{w}(t)\;=\;\operatorname*{arg\,max}_{j:\,\sigma_{j+1}\geq\varepsilon\,\sigma_{1}}\frac{\sigma_{j}}{\textstyle\sum_{i}\sigma_{i}}\cdot\frac{\sigma_{j}}{\sigma_{j+1}},$
where $\varepsilon\ll 1$ floors out negligible eigenvalues. The weight $\sigma_{j}/\sum_{i}\sigma_{i}$ ensures only ratios at eigenvalues carrying substantial variance can determine $k^{*}$. In the grokking experiments (Table 9), the weighted definition recovers $k^{*}=1$ in 9/12 runs (vs. 4/12 unweighted). In the extreme aspect ratio regime the two definitions coincide.
@@ 6.3 Null Distribution of the Maximum Ratio
Under the null hypothesis $H_{0}$: all signal strengths are equal ($d_{1}=d_{2}=\cdots=d_{W}$), the ordered eigenvalues of the $W\times W$ Gram matrix follow a Laguerre ensemble [7, 8]. The distribution of the maximum consecutive ratio $R_{\max}=\max_{j}\lambda_{j}/\lambda_{j+1}$ satisfies, for $W$ fixed and $p\to\infty$:
@@ 9 Subspace Stability: The Davis–Kahan Connection
The key observation is that the Davis–Kahan $\sin\Theta$ theorem [9] (see also Stewart [11] and Kato [10]) does not care what lies below the gap—noise or subdominant signal. It only cares about the size of the gap.
@@ 10.1 Dyson Eigenvalue Dynamics
Corollary 10.2 (Eigenvalue Repulsion [12]).
@@ 12.3 Simplified Signal Flow
Weight decay must sit between the NTK eigenvalues of the generalizing and memorization modes. Too small: memorization persists (no gap opens). Too large: generalization is also suppressed. When $\omega=0$: all low-curvature modes persist at $d_{j}\propto 1/\sqrt{h_{j}}\to\infty$, the generalizing direction is buried, and grokking never occurs. This is consistent with the universal finding of 24/24 grokking with $\omega>0$ and 1/24 without in the Gram matrix (Section 27).
@@ 17.1 Why the Largest Gap: The Gap Maximality Principle
The gap maximality principle is confirmed by every phase transition observed in our experiments (Section 21Section 28): all 24/24 Gram-matrix grokking events with weight decay, the $k^{*}$ shifts in GPT-2 and TinyStories, and the Dyck-1/SCAN gap openings all occur at the position of the largest consecutive eigenvalue ratio, never at secondary gaps.
@@ 27 Test 7: Gap Opening and Grokking (Dyck/SCAN)
Proposition 10.7 predicts that gap opening corresponds to capability gain, and Remark 10.8 maps grokking onto a delayed gap opening event within the signal hierarchy. We test this in a setting where the capability gain is unambiguous: the grokking transition [26, 27], where generalization accuracy jumps from near-zero to near-perfect.
@@
We analyze weight-matrix singular values $\sigma_{1}\geq\sigma_{2}\geq\cdots$ of the query projection $W_{Q}$ during training of 2-layer transformers on Dyck-1 language ($d_{\mathrm{model}}=128$, $\sim$}150K parameters) and 6-layer encoder–decoder transformers on SCAN compositional generalization ($d_{\mathrm{model}}=256$, $\sim$}1.5M parameters). Each task is trained with weight decay $\omega\in\{0,1.0\}$ across 3 seeds, giving 12 runs total. The gap ratio $R(t)=\sigma_{1}(t)/\sigma_{2}(t)$ is computed at each checkpoint. Grokking is defined as the first step where test accuracy exceeds 0.95.
@@ Results.
    Controls: Without weight decay, no gap opens and no grokking occurs—6/6 grokking runs with $\omega=1.0$, 0/6 with $\omega=0$ (100% hit rate, 0% false positives).
@@
Figure 4: Grokking as a spectral edge event (Dyck-1). Grokking runs (weight decay $\omega=1.0$, red/solid, 3 seeds) vs. control runs ($\omega=0$, blue/dashed, 3 seeds). Top: singular value ratio $\sigma_{1}/\sigma_{2}$ of $W_{Q}$. With weight decay, the ratio rises dramatically as $\sigma_{2}\to 0$ (gap opens); without weight decay, the ratio stays flat (no gap). Middle: Frobenius norm $\|W_{Q}\|_{F}$. Weight decay compresses the weights; the control runs retain large norms. Bottom: test accuracy. Grokking (accuracy $\to 1$) occurs only in the weight-decay runs, coinciding with the gap opening above. All 6 grokking runs show gap opening; all 6 controls show neither gap opening nor grokking.
Gram-matrix replication (Dyck/SCAN).
To verify the Dyck/SCAN result in the trajectory Gram matrix, we compute the rolling-window eigenvalue gap $g_{23}=\sigma_{2}^{2}-\sigma_{3}^{2}$ from the Gram matrix of flattened attention-weight updates (all $W_{Q}$, $W_{K}$, $W_{V}$, $W_{O}$ across all layers). We use $W=3$ for Dyck (p=131,072p=131{,}072) and $W=5$ for SCAN (p=2,359,296p=2{,}359{,}296), constrained by checkpoint density before grokking.
[…]
Task Seed Grok $g_{23}^{\mathrm{early}}$ $g_{23}^{\mathrm{grok}}$ Decline $k^{*}_{\mathrm{term}}$
Dyck ($\omega{=}1$) 42 600 3.88 0.081 $48.1\times$ 1
Dyck ($\omega{=}1$) 137 1400 11.34 0.240 $47.2\times$ 1
Dyck ($\omega{=}1$) 2024 1000 10.11 2.230 $4.5\times$ 1
SCAN ($\omega{=}1$) 42 3000 27.73 0.555 $49.9\times$ 1
SCAN ($\omega{=}1$) 137 4000 26.59 0.858 $31.0\times$ 1
SCAN ($\omega{=}1$) 2024 3000 8.21 0.166 $49.5\times$ 1
Dyck ($\omega{=}0$) all gradual decay, no sharp transition 1
SCAN ($\omega{=}0$) all gradual decay, no sharp transition 1
The Gram-matrix $g_{23}$ decline (mean $33\times$ for Dyck, $43\times$ for SCAN) confirms that the update trajectory compresses onto a rank-1 subspace at the grokking transition. Control runs also show $g_{23}$ decline over the full trajectory, but this reflects gradual settling into a memorization minimum rather than a sharp phase transition; the grokking runs concentrate their decline around the grokking step.
This provides a dual view of the same event: in the weight matrix, the gap opens ($\sigma_{1}/\sigma_{2}\to 10^{3}$); in the update Gram matrix, the sub-leading gap closes ($g_{23}\to 0$). Both are consistent: weight decay drives subdominant update directions to zero, simultaneously opening the weight-matrix gap and collapsing the update spectrum.
@@
Replication with the trajectory Gram matrix.
To verify that the $g_{23}^{(W)}$ decline is not an artifact of weight-matrix eigenvalues, we replicate the analysis using the rolling-window Gram matrix $\bm{G}=\bm{X}\bm{X}^{\top}$ of flattened attention-weight updates ($W=10$) on a different architecture (2-layer transformer, $d=128$, 4 heads). Here $g_{23}^{(G)}=\sigma_{2}^{2}-\sigma_{3}^{2}$ denotes the sub-leading gap of the Gram eigenvalues (to distinguish from $g_{23}^{(W)}$ above). Table 9 shows the results for 4 operations $\times$ 3 seeds:
Table 9: Gram-matrix sub-leading gap decline for modular arithmetic (2-layer transformer, d=128d{=}128). The gap $g_{23}=\sigma_{2}^{2}-\sigma_{3}^{2}$ of the rolling-window Gram matrix declines in 12/12 grokking runs (mean $40.2\times$) and 1/12 controls (the anomalous late-grokking $x^{2}{+}y^{2}$ run). Weighted $k^{*}$ reaches 1 in 9/12 runs.
add 42 3100 14.70 0.63 $23.3\times$ 1
add 137 3000 14.74 0.93 $15.8\times$ 1
add 2024 2500 14.43 0.13 $110.1\times$ 8
sub 42 3300 14.39 0.52 $27.7\times$ 1
sub 137 3300 14.55 0.26 $55.4\times$ 1
sub 2024 3700 14.34 0.67 $21.5\times$ 1
mul 42 2600 14.70 0.17 $86.4\times$ 1
mul 137 3000 14.81 0.54 $27.4\times$ 9
mul 2024 2600 14.64 0.46 $32.2\times$ 9
$x^{2}{+}y^{2}$ 42 2000 18.73 1.25 $15.0\times$ 1
$x^{2}{+}y^{2}$ 137 2500 18.63 0.36 $52.0\times$ 1
$x^{2}{+}y^{2}$ 2024 2100 19.47 1.22 $15.9\times$ 1
The decline is substantially larger ($40.2\times$ vs. $6.6\times$), attributable to the richer spectral structure of the 2-layer model’s Gram matrix compared to the 1-layer model’s weight eigenvalues. The gap ratio $R$ of the Gram matrix cleanly separates the two conditions: $R_{\mathrm{early}}=1.40\pm 0.07$ (WD == 1.0) vs. $2.58\pm 0.54$ (WD == 0.0), with non-overlapping standard deviations. The higher $R$ in the control runs reflects rank-1 concentration of memorization updates—the model grinds on training examples with highly aligned gradients—while weight decay spreads updates across spectral modes during circuit formation.
@@
[…]
  • Modular arithmetic: sub-leading gap closes before grokking, triggering a $k^{*}$ shift from position 2–3 to position 1. This is a gap closing event (at the old $k^{*}$) followed by a gap opening event (at the new $k^{*}=1$). The Gram-matrix replication confirms the same pattern with even larger effect ($40\times$, Table 9).
Both mechanisms are predicted by the gap dynamics (Theorem 14.2): gap closing destabilizes the subspace, and gap opening at a new position stabilizes the generalizing direction. The weight decay $\to$ grokking correspondence holds in the Gram matrix: 24/24 grok with $\omega>0$ and 1/24 without (6 Dyck/SCAN + 12 single-task + 6 multi-task modular arithmetic grokking runs, with the single exception being an anomalous late-grokking $x^{2}{+}y^{2}$ control).
@@
Weight decay plays a dual role across all experiments: in Dyck/SCAN, it drives subdominant singular values toward zero (closing the Gram-matrix $g_{23}$); in modular arithmetic, it compresses the sub-leading eigenvalue spectrum (closing $g_{23}$), which triggers spectral symmetry-breaking and a $k^{*}$ shift. The correspondence between weight decay and grokking—24/24 in the Gram matrix, with only 1 anomalous control grokking out of 24 total controls—is consistent with weight decay being a key driver by which gap dynamics are associated with the generalization transition.
@@
The original grokking experiments analyze weight-matrix spectra ($\sigma_{j}(W_{Q})$ for Dyck/SCAN, eigenvalues $\lambda_{j}$ for modular arithmetic). The Gram-matrix replications (Tables 7 and 9) confirm that the same $g_{23}$ decline appears in the trajectory Gram spectrum across all three task families: Dyck ($33\times$), SCAN ($43\times$), and modular arithmetic ($40\times$). For Dyck/SCAN, the two views are complementary faces of the same event: weight decay drives subdominant update directions to zero, simultaneously opening the weight-matrix gap and collapsing the Gram sub-leading gap. The Davis–Kahan stability framework (Remark 16.3) applies to both: gap dynamics control subspace stability regardless of the matrix being analyzed.
@@ 29 Summary of Theory–Experiment Match
Gap dynamics (10.7) gap event $\Rightarrow$ capability 24/24 grok, 1/24 ctrl
@@ 40 Connection to the Edge of Stability
Across all models studied ($k^{*}\in\{1,2,3\}$: $k^{*}=1$ for Muon and grokking tasks, $k^{*}=2$33 for AdamW on GPT-2 and TinyStories), the number of simultaneously active modes is small. A plausible explanation is that when $k>k_{\mathrm{crit}}$ modes simultaneously overshoot, nonlinear interactions destabilise the self-correction mechanism of GD, so the system self-organises to keep only a few modes at the edge. This would make $k^{*}$ a property of the optimizer rather than the task or architecture. We do not have a proof of this; it remains an open question.